A ship’s fuel oil (oil density = 800 kg/m3 ) storage tanks contain oil to a depth of 8 m. If the barometer reading is 750 mm Hg. Determine:

i. The gauge pressure on the plate

ii. The load on the plate due to the oil, if the plate is circular with a diameter of 450 mm

iii. The absolute pressure on the plate 


1
Expert's answer
2021-05-30T13:28:40-0400

a)

"\sigma=800\ kg/m^3" is oil density,

"h=8\ m" is depth.

The gauge pressure on the plate:

"P_G=\sigma gh=800\cdot9.8\cdot8=62720\ Pa"


b)

The load on the plate:

"F=P_G\cdot\pi r^2=62720\cdot\cdot3.14\cdot0.45^2=39990.74\ N"


c)

The absolute pressure on the plate:

"P_{abs}=P_G+P_{atm}"

"P_{atm}=750\ mm\ Hg=750\cdot133.32=99990\ Pa"

"P_{abs}=62720+99990=162710\ Pa"